Determine the distance between 2 − i and 3 + i on the complex plane.


1
Expert's answer
2021-12-26T17:44:46-0500
"|(2-i)-(3+i)|=|-1-2i|"

"=\sqrt{(-1)^2+(-2)^2}=\sqrt{5}"

The distance between "2 \u2212 i" and "3 + i" on the complex plane is "\sqrt{5}."